OpenStudy (anonymous):

An open top box with rectangular base is to be constructed. The box is to be at least 2 inches wide and twice as long as it is wide. It is to have a volume of 150 cubic inches. What should the dimensions of the box be if the surface area is to be as small as possible?

OpenStudy (anonymous):

So, length(l)=2*breadth(b)

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

yeah area would =2xy

OpenStudy (anonymous):

Now, u know the formula for VOLUME OF A CUBOID?

OpenStudy (anonymous):

i dont think so

OpenStudy (anonymous):

Volume of a cuboid=l*b*h

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

where do i go from there?

OpenStudy (anonymous):

Also, Volume =150

OpenStudy (anonymous):

given right?

OpenStudy (anonymous):

yes that was given

OpenStudy (anonymous):

im still not sure how to put this together to reach an answer

OpenStudy (anonymous):

OK: VOlume=150 l*b*h=150 l*2l*h=150 h=75/l^2

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

so height is 75?

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

Now, we have, b=2l h=75/l^2

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

@skallan01 ???

OpenStudy (anonymous):

yeah i got it

OpenStudy (anonymous):

Now, S.A of that rectangular box = 2(l+b)*h + l*b

OpenStudy (anonymous):

following?

OpenStudy (anonymous):

ok i think i can take it from there,thank you

OpenStudy (anonymous):

WELCOME